OpenStudy (anonymous):

A regulation basketball has a diameter of 9.4 inches. What is the volume of the ball? A) 19.69 in3 B) 92.53 in3 C) 434.89 in3 D) 3479.14 in3

jhonyy9 (jhonyy9):

- so than 4,7 inches is the radius V=4/3 (pir^3) =4/3 *(103,823) =138,43*pi=470,662 in^3 - so than C. will be right from your answer

OpenStudy (kira_yamato):

Let's find radius r cm first r = 9.4/2 = 4.7 V = 4/3 π r^3 V = 4/3 π (4.7^3) = 434.89 in3 ==> C